Opened 8 years ago

Closed 3 years ago

#405 closed flight processing (fixed)

GB11_03, flight day 169/2011, Tingwall

Reported by: knpa Owned by: knpa
Priority: whenever Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/uk/GB11_03-2011_169_Tingwall

Data arrived from ARSF via network transfer on 22/06/2011

GB11_03 Scientific objective:
Archaeological investigation of Medieval North Europe using LIDAR survey.

GB11_03 Priority: pilot

PI: Alexandra Sanmark

Sensors:

Camera (15/08/11)
LiDAR (6/9/11)

Change History (25)

comment:1 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Lidar processing currently blocked because of the lack of calibration flight.

Starting camera processing.

comment:2 Changed 8 years ago by knpa

Created Tiffs (on windows box).

comment:3 Changed 8 years ago by knpa

Tiffs 1 to 14 are not on any flightline and will be removed.

comment:4 Changed 8 years ago by knpa

No underexposed/overexposed images. No images of bay doors.

comment:5 Changed 8 years ago by knpa

IPAS navigatio processing complete (camera .sol file created too)

comment:6 Changed 8 years ago by knpa

Creating RCD delivery. Having trouble with the ReadMe: as there are no eagle/hawk lines, the script cannot insert the thumbnails, sorted by line, into the pdf; having to do it manually.

comment:7 Changed 8 years ago by knpa

RCD ready for delivery check at GB11_03-169-camera-20110805

comment:8 Changed 8 years ago by jaho

Camera delivery check:

Kml file was named incorrectly. Renamed to GB11_03-2011_169.kml.
Kml file is not listed in the ReadMe.

No other errors found.

comment:9 Changed 8 years ago by adbe

Awaiting OS Vectors

comment:10 Changed 8 years ago by knpa

Added kml file to Readme contents.

Preparing disk for dispatch.

comment:11 Changed 8 years ago by knpa

Delivered RCD today (15/08/11) to Alexandra Sanmark at the UHI institute.

comment:12 Changed 8 years ago by knpa

  • Description modified (diff)

comment:13 Changed 8 years ago by adbe

Lidar Blocked, awaiting OS vectors

comment:14 Changed 8 years ago by adbe

Checked against vectors

comment:15 Changed 8 years ago by adbe

Ready for delivery check

comment:16 Changed 8 years ago by emca

LiDAR delivery check complete. No issues with data.

Ready for delivery.

comment:17 Changed 8 years ago by emca

LiDAR delivery:
DEM has SRTM in the title but there is no SRTM data patched to either DEM in the delivery folder. Can you re-create these DEMS with the SRTM or other available DEM data added, or extrapolate if no external DEM data available.


comment:18 Changed 8 years ago by adbe

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

The LiDAR data has been sent to Alexandra Sanmark on 6/9/11

comment:19 Changed 7 years ago by knpa

Beginning archiving.

comment:20 Changed 7 years ago by knpa

No lidar intensity screenshot with vectors.

comment:21 Changed 7 years ago by knpa

Above comment is not exactly accurate, the vectors are present on the intensity image but there is no intensity image without the vectors.

comment:22 Changed 6 years ago by emca

Email from Joris Coolen @ Uni Vienna who is working on this data. He has found height offset between LiDAR and GCPs.

Processed the base station data using PPP giving:
60 08 17.98782; -1 11 01.39945; 130.446

Reprocessed the navigation data.

comment:23 Changed 6 years ago by emca

Reprocessed the lidar data and copied the classification over from previous delivery.

Now creating a new delivery.

comment:24 Changed 4 years ago by dac

Starting uploading reprocessed LiDAR data to NEODC.

comment:25 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Reprocessed LiDAR available from NEODC. Marking as closed.

Note: See TracTickets for help on using tickets.